40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Sep 24 17:12:54 2011, Koji, Update, LSC, Realignment of REFL / Some 3f PRMI locking / Recycling Gain 
    Reply  Sat Sep 24 20:14:36 2011, Koji, Update, LSC, Rough estimation of the PR gain 
    Reply  Mon Sep 26 15:15:45 2011, Anamaria, Update, LSC, Realignment of REFL / Some 3f PRMI locking / Recycling Gain 
Message ID: 5545     Entry time: Mon Sep 26 15:15:45 2011     In reply to: 5539
Author: Anamaria 
Type: Update 
Category: LSC 
Subject: Realignment of REFL / Some 3f PRMI locking / Recycling Gain 

A few comments on REFL table alignment and REFL165.

Last time we realigned the table was after the PZT work by Koji/Kiwamu; we made sure that the beam was going through optics satisfactorily and that we were reading reasonable numbers. I did use primarily a viewer to align onto PD, after which we used the voltage reading to center better around that spot. As desired, I could not see the beam once it was centered on the PD. I never touched the PBS unfortunately, so I never noticed it was not fixed. Sad.

I am very surprised to hear the reading from REFL165, since I was reading around 400mV from it a few days before. Something strange happened in the mean time. I hope not when I was plugging and unplugging at the power rack for the POY work. But I would not have needed to touch REFL165. Those cables should get some strain relief at the rack, by the way.

I thought about it, and I must admit that after we centered camera on REFL (paired with an alignment), we did not check the beam path later, even after we saw that the REFL beam had moved. We only did a quick by-viewer check that the beams were not off of the PDs.

Quote:

[Koji Suresh]

- The REFL path has been thoroughly aligned
Many optics had the spots not on the middle of the optic, including the PBS whose post was not fixed on the post holder.
We aligned the optical paths, the RF PDs, and the CCD. The alignment of the PD required the use of the IR viewer.
One should not trust the DC output as a reference of the PD alignment as it is not enough sensitive to the clipping.

We aligned the optical paths again after the reasonable alignment of PRM is established with the interferometer.
"Next time when you see REFL spot is not at the center of the camera, think what is moved!"

- The REFL165 PD is disconnected from the power supply
I found that the REFL165 PD is producing 7.5V output at the DC monitor no matter how the beam is blocked.
As I could not recover this issue by swapping the power connector at the LSC rack, I disconnected the cable
at the RFL165 PD side. I need to go through the PD power supply circuit next week.

 

ELOG V3.1.3-